### v1

# Quantifying the generation of negatively charged boron vacancies in He-ion irradiated hexagonal boron nitride

Abstract
Hexagonal boron nitride (hBN) hosts luminescent defects possessing spin qualities compatible with quantum sensing protocols at room temperature. Vacancies, in particular, are readily obtained via exposure to high-energy ion beams. While the defect creation mechanism via such irradiation is well understood, the occurrence rate of optically active negatively charged vacancies (V–B) is an open question. In this work, we exploit focused helium ions to systematically generate optically active vacancy defects in hBN flakes at varying density. By comparing the density-dependent spin splitting measured by magnetic resonance to calculations based on a microscopic charge model, in which we introduce a correction term due to a constant background charge, we are able to quantify the number of V–B defects generated by the ion irradiation. We find that only a small fraction (0.2%) of all vacancies is in the optically active, negatively charged state. Our results provide a protocol for measuring the generation efficiency of V–B, which is necessary for understanding and optimizing luminescent centers in hBN.
I Introduction
Luminescent centers in hBN emerged as bright and stable room temperature quantum emitters [1, 2], rivaling established platforms in quantum communication and sensing applications [3, 4, 5]. Emitters in hBN exhibit robust photoluminescence (PL) across a broad spectral range from the near-infrared [6] through the visible [1, 7] up to the blue [8, 9, 10, 11, 12] and UV [13] regions, as well as magnetic response both as ensemble spin-1 systems [6] and as single spin-1 [14, 15] or spin-1/2 [16] emitters. Among the most studied types of luminescent centers in hBN is the negatively charged boron vacancy (V–B), which consists of an electron localized at the position of the missing atom in the hexagonal lattice. These defects are typically addressed in PL ensemble measurements, where they exhibit a very broad spectrum centered around 800 nm [6, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Importantly, V–B defects feature a spin-triplet ground state with a zero-field splitting of around 3.5 GHz [6], a corresponding triplet excited state with a splitting of around 2.1 GHz [25, 31, 32, 33], and a metastable state acting as a non-radiative channel with a very short lifetime [29, 30]. The spin state can be optically read-out and coherently manipulated through optically detected magnetic resonance (ODMR) measurement schemes [6, 28, 34, 35, 30]. The microwave resonances are sensitive to external magnetic fields [6, 25], which can be applied for quantum sensing schemes [36, 37, 38, 18], similar to nitrogen vacancy-centers in diamond [39, 40]. Because of their potential for applications in quantum technologies, the optical and magnetic properties of V–B defects have been a subject of intensive research, including their optical response at room [6, 25, 41, 28, 30] and cryogenic [20, 41] temperatures, as well as ODMR [6, 20, 31, 33], electron paramagnetic resonance (EPR) [6, 42], and spin-coherence measurements [4, 43, 44, 45]. Furthermore, several studies discuss various sample preparation and optimization protocols, such as irradiation sputtering [22], ion beam fluence and energy tuning [21, 41], high-temperature annealing [46, 23, 24], water exposure [22], as well as the integration of defects into photonic circuits, such as nanocavities [47, 48], plasmonic- [19] and metasurfaces [49].
For photonic integration reliable, quantitative and deterministic methods for defect generation are desirable. In this context, different methods were successfully applied to create V–B in hBN, which include focused ion beam irradiation using H [17], He [46, 28, 41, 19, 23, 21, 50], N [30, 21, 20], Ga [6, 31], C [21, 25], Xe [20] and Ar [21, 20], neutron irradiation [27, 6, 38, 51, 52], femtosecond laser writing [53], and electron beam irradiation [42, 54]. In particular, several works discussed the usage of a focused He-ion beam source to effectively generate V–B defects in hBN [28, 41, 19, 23, 21, 50]. Focused He-ion beam exposure offers a nominal lateral defect positioning accuracy below 10 nm in very thin films [55], while preserving the local crystal structure of hBN at adequate doses. All such nanofabrication measures are crucial factors for an efficient integration of such luminescent centers in future photonic [56] and quantum imaging devices [18]. In this context, several works [41, 28, 23] already provided initial insights onto the He-ion generation process and the optical response of the generated ensemble emitters, but a systematic understanding, in particular a reliable quantification of defect density and its impact on the magnetic and optical properties of the generated emitters, is still lacking.
In the present work, we use a helium ion microscope (HIM) to irradiate hBN flakes, deposited on thick graphite flakes, to generate V–B defects (Figure 1 a). The thick graphite acts as both a mechanical and optical protection layer, reducing the impact of back-scattered ions from the SiO2 substrate and suppressing unwanted background luminescence from the irradiation damage in the substrate. We systematically perform spatially- and time-resolved PL as well as ODMR measurements on defect ensembles for varying irradiation fluences, confirming the optical and magnetic activity of the generated emitters. We show that the focused He-ion beam technique can site-selectively produce luminescent and background-free emitters, in good agreement with the existing literature [6, 28, 41, 19, 23, 21, 50]. Furthermore, we apply a microscopic charge model [44, 52] to describe the defect density-dependent shift of the ODMR resonance frequencies. By introducing a background charge correction, we can quantify from the experimental data the density of the charged defects generated by the ion exposure. By comparison to molecular dynamics (MD) simulations, which estimate the number of total vacancies generated, we are able to quantify an overall efficiency of the charged defect creation process based on the focused He-ion beam irradiation. Our work provides a systematic study of defect production efficiency by comparing different irradiation fluences on homogeneous and uniform samples, offering a benchmark for future process optimization and device implementation.
II Methods
Thin layers of hBN and graphite are exfoliated using adhesive tape on Si/SiO2 substrates. Subsequently, hBN/graphite heterostructures are assembled with a hot-pick-up technique [57] and cleaned by immersion in chloroform. The individual flake thickness is assessed with atomic force microscopy (AFM). To generate luminescent defects in hBN, we use a focused He-ion beam (ion energy of ) to pattern an array of circular patches with diameter and 2 µm pitch with ion fluences as defined in Table 1. Optical characterization is performed in a custom confocal microscope featuring a 532-nm laser and a grating spectrometer. Continuous-wave (cw) ODMR measurements with the PL integrated from 650 nm to 850 nm are performed with 532 nm laser excitation and a microwave excitation at 3 GHz – 4 GHz. Molecular dynamics (MD) simulations are carried out using the LAMMPS package [58], to extract the statistics of defects produced by the impacts of He ions onto the hBN. More details about the methods can be found in Appendix A and the motivation for the simulation setup is discussed in Refs. [59, 60, 61].
III Results
| # | Dose | Fluence | V–B density | V density | V–B/V | |
| 1 | 5000 | 0.0343(38) | 0.0059(51) | 7.67(9) | 0.2 % | |
| 2 | 1000 | 0.0296(36) | 0.0012(49) | 1.53(2) | 0.2 % | |
| 3 | 500 | 0.0291(40) | 0.0006(52) | 0.767(9) | 0.2 % | |
| 4 | 200 | 0.0286(55) | 0.0002(65) | 0.307(4) | 0.1 % | |
| 5 | 100 | - | - | 0.153(2) | - | |
| 6 | 25 | - | - | 0.0384(5) | - | |
| 7 | 10 | - | - | 0.0153(2) | - |
III.1 Optical characterization of V–B centers
Figure 1 b depicts a typical hBN flake (80 nm thick) on graphite supported on Si/SiO2. The sample is exposed with arrays of circular patches of diameter and pitch. In those circular patches, the ion fluence is varied from to (see Table 1). To detect the defect emission, the PL is filtered from 650 nm to 850 nm and recorded with a single-photon counting device. The PL mapping after He-ion irradiation (Figure 1 c) clearly resolves the ion exposed spots by their increased luminescence (Figure A.2 a) as well as other bright spots, including folds and edges. A logarithmic scale is used for the color bar to visualize a larger dynamic range.
Figure 1 d shows representative PL spectra recorded on the exposed spots for different ion fluences. The most prominent luminescence feature is a broad peak () centered around (), whose intensity increases monotonously with increasing He-ion exposure (Figure A.5 a). The PL can be saturated at large powers, about 5 mW for spots with (Figure A.5 b), which is indicative of emission from a finite number of localized defect centers [21, 28, 29, 30]. For comparison, the gray spectrum in Figure 1 d was recorded in a location not exposed to He-ions, i.e. on pristine hBN, where this spectral feature is absent. The peak position and spectral shape of the ion-induced luminescence feature are consistent with other studies that demonstrated emission from V–B [6, 17, 18, 19, 20, 21, 22, 23, 62].
The large broadening, which obscures the observation of a zero-phonon line, is generally attributed to dominant phonon-assisted emission processes from an ensemble of emitters [6, 4, 63]. Cavity-enhanced experiments put the zero-phonon line of V–B at around 770 nm [48], whereas simulations show that phonon-independent luminescence is extremely weak [63]. It should be noted that making a quantitative comparison of the line shape for such a broad luminescence feature across different studies is often challenging, unless meticulous calibration of the overall spectral efficiency of the optics used is performed. This accounts for slight variations between this work and other works [6, 64, 28, 21, 30]. At the highest fluence (, dark purple spectrum in Figure 1 d), a background luminescence emerges across the full spectral range. This increased background correlates with the occurrence of an additional Raman mode at , which has previously been attributed to ion-beam induced damage in hBN [41, 28, 65] and which is absent in the pristine control spectrum (inset of Figure 1 d, see Figure A.1 the for full Raman spectrum). The background luminescence and defect-activated Raman mode may hint towards the onset of ion-induced amorphization of the hBN lattice [28] or the prevalent creation of larger defect clusters [65]. Consequently, up to the second highest fluence of , we can assume that luminescent defects are generated in a rather dilute limit.
Consistently, we find that the PL lifetime does not display any clear change with irradiation fluence (or equivalently the induced defect density) except for the highest irradiation fluence, where the lifetime drops significantly from about to (Figure 2). The lifetimes are in line with values reported by other studies [6, 25, 31, 49, 30]. To fit the experimental data, a double-exponential model is applied, in which the first component captures the lifetime at longer time scales and the second component addresses the fine details at short time scales. The fit includes a convolution with the independently measured instrument response function. The time traces are shown in Figure 2 exemplarily for (red curve) and (blue curve). Table 2 in Appendix A.2 summarizes the extracted time constants (see Figure A.3 for the full lifetime data). Note that for optical emission with low quantum yield, as is the case for V–B, the measured lifetimes in time-resolved PL reflect a combination of the radiative and non-radiative rates. Therefore, the shortening of the lifetime with the fluence indicates an increased non-radiative recombination, which is likely due to the significant damage imparted onto the hBN lattice at the highest fluence. So far, studies reported tuning of the lifetime via temperature [31] or by exploiting spin-dependent non-radiative recombination using specific measurement pulse sequences [25, 30], where first long optical pulses initialize the spin polarization in the state and microwave -pulses rotate the spin polarization in one of the states ) thus shortening the overall lifetime. In our case we are therefore measuring the PL decay dynamics of a population in which both the radiative and non-radiative recombination mechanisms play a role [30].
III.2 Optically detected magnetic resonance and microscopic charge model
The V–B defect is considered to be one of the few hBN defects whose atomic structure and level scheme are well known [6, 62, 64, 25, 31, 44, 52, 51]. Figure 3 e shows the level structure. The term indicates the projection of electron spin along the -axis of the hBN crystal, considered the quantization axis of the system [52]. The system exhibits a spin triplet ground state (GS) , with a so-called axial zero-field splitting around separating the and spin sublevels, and a corresponding spin triplet excited state (ES) , which is split between the spin sublevels as well [62, 64, 30]. Transitions between the ground and excited states are optically addressable under green laser excitation (typical wavelength 532 nm). Additionally, a manifold of metastable singlet states (MS) is present, which can be populated from the excited state via intersystem crossing (ISC). The phonon-assisted radiative emission from ES to GS is rather weak since it is forbidden by symmetry [30], thus these emitters display a low quantum efficiency. On the other side, recent studies found that the MS has a rather short lifetime (on the order of tens of nanoseconds) [29, 30]. Therefore, the defect can be pumped efficiently, and its PL can be detected at sufficient intensity despite the very low quantum efficiency [30].
The ODMR spectra are measured at zero external magnetic field and on four different ion fluences (all on the same sample of Figure 1 c) corresponding to four emitter densities (Figure 3 a). All ODMR spectra display two characteristic resonances. We verified that their splitting increases in frequency under a static external magnetic field [6, 25] (Figure A.6). The slight difference in the overall magnitude of the contrast is due to a varying coupling efficiency of different positions in the sample to the nearby microwave antenna. To extract the splitting between the resonances, we fit the data with two exponentially modified Gaussians, which account for an asymmetric broadening of the resonances (see Figure A.7 b and the discussion in Appendix A.5). Figure 3 b shows the resonance positions and , extracted as the maxima of the individual resonance distribution, as a function of the irradiation fluence as well as their midpoint defined as .
The splitting between the resonances increases with increasing ion fluence, or equivalently defect density (Figure 3 c). In literature, such splitting is attributed to the interaction of the V–B electronic spin with local fluctuations of the electric environment [44, 51, 52]. By contrast, variations of the mean are related to local strain fields [52] and temperature [66]. By modeling the magnitude of the splitting, it is possible to estimate the local field strength and, therefore, the local charge density surrounding the emitters. Following the approach in Refs. 44, 51, 52, we apply a microscopic charge model to simulate the level splitting. Briefly, we consider level splitting due to a net electric field, which is generated by a discrete and random charge distribution in the vicinity of the V–B site. Importantly, the model still assumes overall charge neutrality, i.e. positive and negative point charges are placed around the V–B in equal numbers (see Appendix A.5 for details). The remaining net electric field arises solely from the locally net sum of the field components. To date, the microscopic origin of such a charge environment is not fully understood. A common assumption is that the charge fluctuations simply correspond to the number of charged V–B defects, which, as we discuss in the following, is not fully consistent with our experimental results.
Figure 3 d depicts the simulation results for a selected range of average charge densities, in which we fit the experimentally measured values for different ion fluences (orange data points) to the calculated values of and as a function of charge density (solid lines). In the fitting procedure, we match only the splitting rather than the absolute values of and , such that we eliminate strain effects affecting the mean [52] (see Figure A.7 for full data set and details of fitting procedure). The error bars take into account both the experimental uncertainty, i.e. the fitting error of the peak splitting (cf. Figure 3 c), and the theoretical uncertainty, i.e. the statistical spread of the eigenvalues in the simulation (cf. Figure A.7 c).
Generally, the extracted charge densities are in line with other reports [44, 51, 52]. This finding is consistent with the fact that we measure similar magnitudes of the splitting and incorporate the value of the out-of-plane susceptibility determined in Ref. 44 as the main parameter in the model. However, unlike Ref. 52, we do not observe a proportional scaling of charge density and PL intensity with ion fluence, whereby the latter can be taken as a good estimate of the relative number of optically active defects for a constant lifetime (cf. Figure 2) and for an excitation power below the saturation threshold (Figure A.5). Specifically, the PL intensity of V–B emission increases by a factor of 10 (Figure A.5) as the ion fluence is varied over more than one order of magnitude from to while, at the same time, the average charge density extracted from the model increases only slightly from to . Therefore, as a possible extension of the model, we propose that the overall charge density should include a constant positive and negative background charge, which is present before ion irradiation. The origin of such a background charge may potentially be related to native defects, such as interstitials, impurities, or vacancies present after the growth. Charge transfer processes between such defects can then lead to local charge accumulation, while preserving overall charge neutrality. Such a background charge would lead to a finite level splitting already in the limit of small V–B defect density, as we clearly observe for (Figure 3 a). With these assumptions, we determine a background charge of . From that, we can find the corrected values for the V–B density as given in Table 1.
III.3 Molecular dynamics simulation of vacancy creation
In a next step, we correlate the experimentally found production yield to molecular dynamics (MD) simulations of the ion matter interaction. We model the 30 keV He-ions impinging onto a -thick hBN sheet. The calculations are performed for a reference fluence (see Appendix A.6) and incidence angle. Contrary to the binary collision approximation (BCA), used in common SRIM (stopping range of ions in matter) simulations [21, 41, 19, 28, 17], MD simulations take into account the molecular structure of the target materials (an example is shown in Figure 4 c), which yields more precise estimates of the defect production. Figure 4 a depicts the depth profile of nitrogen (blue) and boron (red) vacancies at the reference fluence corresponding to . On average, we find densities of and for nitrogen and boron, respectively. Figure 4 b compares the MD approach to simulations based on BCA. For the latter we find on average higher defect densities, which agrees with the general notion that these simplified methods tend to overestimate the defect creation.
We thus compare the results from the MD simulations and from the microscopic charge model to estimate the production efficiency of the irradiation process. The results are shown in the last column of Table 1. Overall, our analysis shows that the production efficiency is rather low at about 0.2 %, implying that although with larger implantation fluences more vacancies are created, a large fraction of vacancies does not get promoted to V–B and remains optically inactive [67, 68, 44]. It should be pointed out that our simulations do not account for in situ annealing of defects through annihilation of vacancies with interstitials, which are mobile at room temperature [69, 70],as it happens at a macroscopically long time scale. The annealing should decrease the number of vacancies, independently of their charge state. Our values for the defect yield constitute a lower bound, since we included the potential effect of a constant background charge. In contrast, Ref. 44 reported a larger negatively charged V–B fraction of % for a fluence , which is comparable to our lowest irradiation fluence for which ODMR was measured. However, we note that the latter study concludes that the defect yield decreases steeply with ion fluence. With our background correction, the derived defect yield is constant within the experimental uncertainty. In other words, we tentatively state that the microscopic charge model without background correction tends to significantly overestimate the V–B density at low ion fluences.
IV Conclusions
In summary, we quantified the defect creation within homogeneous hBN films using a helium ion microscope to site-selectively pattern of V–B defect ensembles. We analyzed the optical properties of the emitters as a function of ion fluence over more than two orders of magnitude. The spatially precise generation allowed us to study the magnetic response probed by ODMR contrast measurements in a consistent and systematic way. We simulated the level splitting obtained experimentally from the ODMR spectra employing a microscopic charge model, where the spatially random distribution of local charges results in a level splitting in ODMR. As a main finding, we observe a discrepancy between the scaling of the charge density, or equivalently the level splitting in ODMR, and the scaling of the defect PL, or equivalently the density of optically active defects, as function of ion fluence. Unlike other studies, which attributed the charge density in the microscopic model solely to V–B defects, we propose to include a background density in the model. With this modification, we were able to estimate the defect creation yield and compare it to molecular dynamics simulations. Our results systematically quantify the He-ion based creation of V–B defects in hBN. On the one hand, further developments may involve implementing new theoretical models with the aim of addressing the fundamental question of the origin of the V–B charge [71], such as the nature of the charged state (dynamical or static) and the role of the local defect environment, as has been done for N centers [72]. On the other hand, spatially precise creation of such emitters as demonstrated here may enable their integration in gated, light-emitting hetero-stacks, in which a tunneling current can be coupled to the in-gap defect states by means of graphene electrodes [56, 73].
Acknowledgements.
N. S. thanks the Novo Nordisk Foundation NERD Programme (project QuDec NNF23OC0082957). Work at the Technical University of Munich was supported through the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the Munich Center for Quantum Science and Technology No. (MCQST)-EXC-2111-390814868. C.K., L.G., J.F. and A.W.H. acknowledge funding through the International Graduate School of Science and Engineering (TUM-IGSSE, project BrightQuanDTUM). N.S. and M.W. acknowledge the support from the Independent Research Fund Denmark, Natural Sciences (project no. 0135-00403B) and from the Danish National Research Foundation through NanoPhoton - Center for Nanophotonics (project no. DNRF147). A.H. and I.B. acknowledge the NNF Biomag project (NNF21OC0066526). A.V.K. acknowledges funding from the German Research Foundation (DFG), projects KR 4866/9-1, and the collaborative research center “Chemistry of Synthetic 2D Materials” CRC-1415-417590517. Generous CPU time grants from the Technical University of Dresden computing cluster (TAURUS) and Gauss Centre for Supercomputing e.V., Supercomputer HAWK at Höchstleistungsrechenzentrum Stuttgart are greatly appreciated.Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions
N.S., A.H., A.W.H and C. K. conceived the project. N.S., C.K., A.H., M.W. and A.W.H. supervised the project. A.C. fabricated the sample. J.F. and A.C. performed the HIM irradiation. L.G. performed the AFM measurements. A.C. performed PL, Raman and lifetime measurements. I.B. and A.C performed PL and ODMR measurements. A.C. performed the microscopic charge model simulations. S.K. and A.V.K. performed the MD simulations. J.A. and A.B.M. performed measurements on further samples.
Data Availability
The data that support the findings of this study are available from the corresponding authors upon reasonable request.
Appendix A Methods
A.1 Sample fabrication
Thin layers of hBN (HQ Graphene) and graphite are exfoliated using adhesive tape (Nitto blue tape for hBN and Scotch Magic Tape for graphite) on Si/SiO2 substrates (oxide thickness 90 nm, cleaned by sonication in acetone and subsequently isopropanol). For graphite exfoliation, we pre-heat the substrates, which were beforehand treated in a mild oxygen plasma for 10 minutes. Suitable hBN and graphite flakes are identified by optical microscopy based on size, contrast, and cleanliness. To assemble hBN/graphite heterostructures, we use a hot-pick-up technique based on poly-carbonate (PC)/polydimethylsiloxane (PDMS) stamps. A thin film of PC is placed on a droplet of PDMS on a glass plate. The stamp is brought in contact with the target hBN flake, and the substrate is heated to to pick-up the flake. The hBN is subsequently released on the target graphite flake by melting the PC film in contact with the target substrate at . The samples are cleaned by immersion in chloroform for few minutes. Finally, mild oxygen plasma cleaning () is performed to remove any remaining polymer residues from the sample surface. The cleanliness and individual flake thickness of the prepared heterostructures are characterized by standard atomic force microscopy (AFM, neaSNOM Attocube Systems AG) in tapping mode.
A.2 Optical Characterization
| # | (ns) | (ns) |
| 1 | 0.774(1) | 0.031(1) |
| 2 | 0.996(3) | 0.070(4) |
| 3 | 1.073(5) | 0.138(8) |
| 4 | 1.028(3) | 0.031(2) |
| 5 | 1.113(9) | 0.017(3) |
Optical properties are characterized by diffraction-limited confocal scanning photoluminescence (PL) mapping, Raman microscopy, and time-resolved photoluminescence (TRPL) under ambient conditions. PL measurements are performed on a custom confocal microscope featuring a 532-nm laser (Cobolt, Hübner GmbH). The objective is a Nikon TU Plan Fluor /0.9 NA EPI D. The signal is collected by a spectrometer (Andor-Solis SR-303i, 150 grooves/mm grating with blaze at 800 nm) equipped with a CCD (Andor Newton). Before the He-ion beam irradiation described below, we verify that the selected hBN flakes contain low background luminescence in the spectral range where emission from V–B is expected ( – ) as well as that only a few intrinsic quantum emitters, with typical emission lines in the visible, are present. Raman spectroscopy is carried out in a commercial Raman microscope (WiTec Alpha, 532 nm laser, cw-excitation, 300 grooves/mm grating with blaze at 500 nm, Zeiss EC Epiplan-Neofluar objective lens / 0.9 NA).
Time-resolved PL measurements (Figure A.3) are performed on the same setup as for PL experiments. For the excitation, we use a PicoQuant 515 nm laser diode in pulsed mode (200 µW of power, 10 MHz repetition rate), in combination with an avalanche photodiode (MPD) and a Time Tagger 20 (Swabian Instruments). The width of the instrument response function (measured on a reflective substrate) is around 250 ps. For fitting, we use the Python package Lifefit [74].
A.3 Helium-ion irradiation
To generate luminescent defects in the hBN films, we use a focused He-ion beam in a helium-ion microscope (Zeiss Orion). As irradiation patterns, we choose arrays of circular patches with diameter and 2 µm pitch. The former is on the order of the typical spot size in the confocal microscope, such that He-irradiated spots appear point-like in the PL images (Figure A.2), and the latter is much larger, so that the exposure spot can be well separated laterally. We use an acceleration voltage of , a beam spacing of 5 nm in x- and y-directions, a beam aperture of , dwell time of and a beam current of . The values of the irradiation fluence are set in on the HIM (see Table 1 in the main text for the fluences employed). To minimize unwanted irradiations, we use a large field of view and low imaging fluence for locating the heterostructures. For hBN placed directly on Si/SiO2, we find significant background PL after ion exposure. The PL originates from both the Si/SiO2 substrate and the hBN (Figure A.4). To avoid such background PL, all the measurements in the main manuscript are conducted on hBN/graphite/SiO2/Si heterostructures, resulting in spectra virtually free from unwanted background luminescence.
The defect PL scales monotonously with ion fluence (Figure A.5 a). Furthermore, for a given defect density, the PL scales sub-linearly with laser power and shows saturation behavior as expected for a finite number of emitters (Figure A.5 b).
A.4 Optically detected magnetic resonance (ODMR)
We perform continuous-wave (CW) ODMR-measurements using a closed-loop coaxial antenna without electrical termination, positioned between the objective and the sample. This configuration allows the emitted light to pass through the loop. The microwave field is generated with an SRS SG384 signal generator and a +40 dBm microwave amplifier. For optical excitation, we use a laser at 532 nm, 6.0 mW with objective lens 60/0.7 NA. With the microwave excitation applied between 3 GHz – 4 GHz, we detect the integrated PL (range 650 nm – 850 nm). For control measurements, a magnetic field is applied through a permanent magnet placed in close proximity to the flake Figure A.6.
A.5 Microscopic charge model
We model the V–B defect as a net negative charge situated at one lattice site surrounded by a number of negative and positive charges. The charges are localized on random sites of the hexagonal lattice and positive and negative charges are assumed to occur in equal numbers, , to preserve overall charge neutrality within a fixed shell volume with radius . We exclude charges on sites which are closer than two times the inter-atomic distance (a= [75]) effectively neglecting multi-vacancies or defect complexes and cutting off strong fields from very close charges that would otherwise lead to divergent behavior.
Such a random charge distribution creates a net electric field on the V–B lattice site, whose in-plane components , perpendicular to the quantization axis, mix the spin sublevels . The mixing leads to new eigenstates with energies separated by a splitting . The out-of-plane field components are not considered because the component of the susceptibility vanishes due to mirror symmetry [44, 52]. The V–B ground state Hamiltonian can then be written as [44, 51, 52]:
with , according to Ref. 44, being the spin-1 operator for V–B electronic spin, being the spin-1 operator for the closest three nuclear spins, and being the hyperfine coupling strength [6]. The nuclear spins are treated as random perturbations not correlated to the state of V–B, and we simply set each as with uniform probability, and average over different spin configurations to extract their contribution [44].
Diagonalizing , we obtain the corresponding eigenvalues as a function of . These are then averaged over 104 possible configurations to take into account the fact that each V–B in the ensemble sees a different charge environment. The full results are shown in Figure A.7 a. Note that, at each charge density, the calculated eigenvalues show a statistical distribution which falls off approximately exponentially with increasing detuning from the zero-field splitting resulting in an asymmetric distribution around the peak maxima (Figure A.7 b).
For such data, calculating simply the mean of the eigenvalues would skew the result towards large detuning and overestimate the corresponding peak splitting. Therefore, we rather quote the position of the peak maxima as the level splitting. To account for such an asymmetric broadening also in the evaluation of the experimental spectra, we use an exponentially modified Gaussian for fitting the ODMR data (see Figure 3 a in the main text). To fit the experimentally derived splitting (in Figure 3 c) to the simulations we apply the procedure outlined in Figure A.7 c. Briefly, we linearly approximate the simulation results in the experimentally relevant range of charge densities. Then, we use this linear fit to match the experimentally measured splitting to a charge density. Lastly, we consider both the variance of the model results, i.e. the error from the linear fit, and the uncertainty of the experiment, i.e. the error bar from fitting the ODMR data, to calculate the error bar of the charge density by Gaussian error propagation.
A.6 Molecular Dynamics (MD) simulations
MD simulations are carried out using the LAMMPS package [58] to extract the statistics of defect creation produced by the interaction of He ions with hBN. The interaction of the ion beam with the hBN structure is modeled via an extended Tersoff potential for boron nitride (BN-ExTeP) [76], with a smooth transition to the Ziegler–Biersack–Littmark (ZBL) potential [77] for small distances. The simulation spans a material thickness of and is performed for a reference fluence, given by a number of 990 ions and a covered surface of 35 µm 35 µm. The MD approach is compared to other Monte Carlo methods, which are the TRIDYN code [78] and the SRIM software [79].

### v3

@@
-Quantifying the generation of negatively charged boron vacancies in He-ion irradiated hexagonal boron nitride
+Quantifying the creation of negatively charged boron vacancies in He-ion irradiated hexagonal boron nitride
 Abstract
@@
 Vacancies, in particular, are readily obtained via exposure to high-energy ion beams.
-While the defect creation mechanism via such irradiation is well understood, the occurrence rate of optically active negatively charged vacancies (V–B) is an open question.
-In this work, we exploit focused helium ions to systematically generate optically active vacancy defects in hBN flakes at varying density.
-By comparing the density-dependent spin splitting measured by magnetic resonance to calculations based on a microscopic charge model, in which we introduce a correction term due to a constant background charge, we are able to quantify the number of V–B defects generated by the ion irradiation.
-We find that only a small fraction (0.2%) of all vacancies is in the optically active, negatively charged state.
-Our results provide a protocol for measuring the generation efficiency of V–B, which is necessary for understanding and optimizing luminescent centers in hBN.
+While the defect creation mechanism via such irradiation is well understood, the occurrence rate of optically active negatively charged vacancies (V–B) is still an open question.
+In this work, we exploit focused helium ions to systematically create optically active vacancy defects in hBN flakes at varying density.
+By comparing the density-dependent spin splitting measured by magnetic resonance to calculations based on a microscopic charge model, in which we introduce a correction term due to a constant background charge, we are able to quantify the number of V–B defects created by the ion irradiation.
+We find a lower bound for the fraction (0.2%) of all vacancies in the optically active, negatively charged state.
+Our results provide a protocol for measuring the creation efficiency of V–B, which is necessary for understanding and optimizing luminescent centers in hBN.
 I Introduction
@@
 The spin state can be optically read-out and coherently manipulated through optically detected magnetic resonance (ODMR) measurement schemes [6, 28, 34, 35, 30].
-The microwave resonances are sensitive to external magnetic fields [6, 25], which can be applied for quantum sensing schemes [36, 37, 38, 18], similar to nitrogen vacancy-centers in diamond [39, 40].
+The spin sub-levels, and thus the two resonances in the ODMR spectrum, are sensitive to external magnetic fields [6, 25], which can be applied for quantum sensing schemes [36, 37, 38, 18], similar to nitrogen vacancy-centers in diamond [39, 40].
 Because of their potential for applications in quantum technologies, the optical and magnetic properties of V–B defects have been a subject of intensive research, including their optical response at room [6, 25, 41, 28, 30] and cryogenic [20, 41] temperatures, as well as ODMR [6, 20, 31, 33], electron paramagnetic resonance (EPR) [6, 42], and spin-coherence measurements [4, 43, 44, 45].
 Furthermore, several studies discuss various sample preparation and optimization protocols, such as irradiation sputtering [22], ion beam fluence and energy tuning [21, 41], high-temperature annealing [46, 23, 24], water exposure [22], as well as the integration of defects into photonic circuits, such as nanocavities [47, 48], plasmonic- [19] and metasurfaces [49].
-For photonic integration reliable, quantitative and deterministic methods for defect generation are desirable.
-In this context, different methods were successfully applied to create V–B in hBN, which include focused ion beam irradiation using H [17], He [46, 28, 41, 19, 23, 21, 50], N [30, 21, 20], Ga [6, 31], C [21, 25], Xe [20] and Ar [21, 20], neutron irradiation [27, 6, 38, 51, 52], femtosecond laser writing [53], and electron beam irradiation [42, 54].
-In particular, several works discussed the usage of a focused He-ion beam source to effectively generate V–B defects in hBN [28, 41, 19, 23, 21, 50].
+For photonic integration reliable, quantitative and deterministic methods for defect creation are desirable.
+In this context, different methods were successfully applied to create V–B in hBN, which include focused ion beam irradiation using H [17], He [46, 28, 41, 19, 23, 21, 50], N [30, 21, 20, 51, 52], Ga [6, 31], C [21, 25, 53], Xe [20] and Ar [21, 20], neutron irradiation [27, 6, 38, 54, 55, 56], femtosecond laser writing [57], and electron beam irradiation [42, 58].
+In particular, several works discussed the usage of a focused He-ion beam source to effectively create V–B defects in hBN [28, 41, 19, 23, 21, 50].
 Focused He-ion beam exposure offers a nominal lateral defect positioning accuracy below 10 nm in very thin films [59], while preserving the local crystal structure of hBN at adequate doses.
 All such nanofabrication measures are crucial factors for an efficient integration of such luminescent centers in future photonic [60] and quantum imaging devices [18].
-In this context, several works [41, 28, 23] already provided initial insights onto the He-ion generation process and the optical response of the generated ensemble emitters, but a systematic understanding, in particular a reliable quantification of defect density and its impact on the magnetic and optical properties of the generated emitters, is still lacking.
-In the present work, we use a helium ion microscope (HIM) to irradiate hBN flakes, deposited on thick graphite flakes, to generate V–B defects (Figure 1 a).
+In this context, several works [41, 28, 23] already provided initial insights onto the He-ion creation process and the optical response of the created ensemble emitters, but a systematic understanding, in particular a reliable quantification of defect density and its impact on the magnetic and optical properties of the created emitters, is still lacking.
+In the present work, we use a helium ion microscope (HIM) to irradiate hBN flakes, deposited on thick graphite flakes, to create V–B defects (Figure 1 a).
 The thick graphite acts as both a mechanical and optical protection layer, reducing the impact of back-scattered ions from the SiO2 substrate and suppressing unwanted background luminescence from the irradiation damage in the substrate.
-We systematically perform spatially- and time-resolved PL as well as ODMR measurements on defect ensembles for varying irradiation fluences, confirming the optical and magnetic activity of the generated emitters.
+We systematically perform spatially- and time-resolved PL as well as ODMR measurements on defect ensembles for varying irradiation fluences, confirming the optical and magnetic activity of the created emitters.
 We show that the focused He-ion beam technique can site-selectively produce luminescent and background-free emitters, in good agreement with the existing literature [6, 28, 41, 19, 23, 21, 50].
 Furthermore, we apply a microscopic charge model [44, 55] to describe the defect density-dependent shift of the ODMR resonance frequencies.
-By introducing a background charge correction, we can quantify from the experimental data the density of the charged defects generated by the ion exposure.
-By comparison to molecular dynamics (MD) simulations, which estimate the number of total vacancies generated, we are able to quantify an overall efficiency of the charged defect creation process based on the focused He-ion beam irradiation.
+By introducing a background charge correction, we can quantify from the experimental data the density of the charged defects created by the ion exposure.
+By comparison to molecular dynamics (MD) simulations, which estimate the number of total vacancies created, we are able to quantify an overall efficiency of the charged defect creation process based on the focused He-ion beam irradiation.
 Our work provides a systematic study of defect production efficiency by comparing different irradiation fluences on homogeneous and uniform samples, offering a benchmark for future process optimization and device implementation.
@@
 The individual flake thickness is assessed with atomic force microscopy (AFM).
-To generate luminescent defects in hBN, we use a focused He-ion beam (ion energy of ) to pattern an array of circular patches with diameter and 2 µm pitch with ion fluences as defined in Table 1.
+To create luminescent defects in hBN, we use a focused He-ion beam (ion energy of ) to pattern an array of circular patches with diameter and 2 µm pitch with ion fluences as defined in Table 1.
 Optical characterization is performed in a custom confocal microscope featuring a 532-nm laser and a grating spectrometer.
-Continuous-wave (cw) ODMR measurements with the PL integrated from 650 nm to 850 nm are performed with 532 nm laser excitation and a microwave excitation at 3 GHz – 4 GHz.
+Continuous-wave (cw) ODMR measurements, with the PL integrated from 650 nm to 850 nm, are performed with 532 nm laser excitation and a microwave excitation at 3 GHz – 4 GHz.
 Molecular dynamics (MD) simulations are carried out using the LAMMPS package [62], to extract the statistics of defects produced by the impacts of He ions onto the hBN.
@@
 In those circular patches, the ion fluence is varied from to (see Table 1).
-To detect the defect emission, the PL is filtered from 650 nm to 850 nm and recorded with a single-photon counting device.
-The PL mapping after He-ion irradiation (Figure 1 c) clearly resolves the ion exposed spots by their increased luminescence (Figure A.2 a) as well as other bright spots, including folds and edges.
+To spatially map the defect emission, we use confocal fluorescence imaging.
+Here, the PL is filtered from 650 nm to 850 nm and recorded with an avalanche photodiode in photon counting mode while the excitation laser is scanned across the sample.
+The fluorescence mapping after He-ion irradiation (Figure 1 c) clearly resolves the ion-exposed spots by their increased luminescence (Figure A.2 a) as well as other bright spots, including folds and edges.
 A logarithmic scale is used for the color bar to visualize a larger dynamic range.
@@
 The most prominent luminescence feature is a broad peak () centered around (), whose intensity increases monotonously with increasing He-ion exposure (Figure A.5 a).
-The PL can be saturated at large powers, about 5 mW for spots with (Figure A.5 b), which is indicative of emission from a finite number of localized defect centers [21, 28, 29, 30].
+Consistent with earlier reports [20, 19, 30], the PL can be saturated at large powers, about 1.37 mW (measured at the entrance aperture of the objective) for spots with (Figure A.5 b), which is indicative of emission from a finite number of localized defect centers [21, 28, 29, 30].
 For comparison, the gray spectrum in Figure 1 d was recorded in a location not exposed to He-ions, i.e.
@@
 At the highest fluence (, dark purple spectrum in Figure 1 d), a background luminescence emerges across the full spectral range.
-This increased background correlates with the occurrence of an additional Raman mode at , which has previously been attributed to ion-beam induced damage in hBN [41, 28, 65] and which is absent in the pristine control spectrum (inset of Figure 1 d, see Figure A.1 the for full Raman spectrum).
+This increased background correlates with the occurrence of an additional Raman mode at , which has previously been attributed to ion-beam induced damage in hBN [41, 28, 69] and which is absent in the pristine control spectrum (inset of Figure 1 d, see Figure A.1 for the full Raman spectrum).
 The background luminescence and defect-activated Raman mode may hint towards the onset of ion-induced amorphization of the hBN lattice [28] or the prevalent creation of larger defect clusters [69].
-Consequently, up to the second highest fluence of , we can assume that luminescent defects are generated in a rather dilute limit.
 Consistently, we find that the PL lifetime does not display any clear change with irradiation fluence (or equivalently the induced defect density) except for the highest irradiation fluence, where the lifetime drops significantly from about to (Figure 2).
@@
 Therefore, the shortening of the lifetime with the fluence indicates an increased non-radiative recombination, which is likely due to the significant damage imparted onto the hBN lattice at the highest fluence.
-So far, studies reported tuning of the lifetime via temperature [31] or by exploiting spin-dependent non-radiative recombination using specific measurement pulse sequences [25, 30], where first long optical pulses initialize the spin polarization in the state and microwave -pulses rotate the spin polarization in one of the states ) thus shortening the overall lifetime.
+So far, studies reported tuning of the lifetime via temperature [31] or by exploiting spin-dependent non-radiative recombination using specific measurement pulse sequences [25, 30] (where first long optical pulses initialize the spin polarization in the state and microwave -pulses rotate the spin polarization in one of the states ), thus shortening the overall lifetime.
 In our case we are therefore measuring the PL decay dynamics of a population in which both the radiative and non-radiative recombination mechanisms play a role [30].
@@
 Figure 3 e shows the level structure.
+The system exhibits a spin triplet ground state (GS) and a corresponding spin triplet excited state (ES) .
+The ground state shows a so-called axial zero-field splitting around separating the and spin sublevels.
 The term indicates the projection of electron spin along the -axis of the hBN crystal, considered the quantization axis of the system [55].
-The system exhibits a spin triplet ground state (GS) , with a so-called axial zero-field splitting around separating the and spin sublevels, and a corresponding spin triplet excited state (ES) , which is split between the spin sublevels as well [62, 64, 30].
+An electric field perpendicular to the quantization axis mixes the sub-levels, resulting in new eigenstates and with a level splitting of (Figure 3 e inset), where denotes the corresponding perpendicular susceptibility.
 Transitions between the ground and excited states are optically addressable under green laser excitation (typical wavelength 532 nm).
 Additionally, a manifold of metastable singlet states (MS) is present, which can be populated from the excited state via intersystem crossing (ISC).
-The phonon-assisted radiative emission from ES to GS is rather weak since it is forbidden by symmetry [30], thus these emitters display a low quantum efficiency.
-On the other side, recent studies found that the MS has a rather short lifetime (on the order of tens of nanoseconds) [29, 30].
-Therefore, the defect can be pumped efficiently, and its PL can be detected at sufficient intensity despite the very low quantum efficiency [30].
+The phonon-assisted radiative emission from excited state to ground state is rather weak since it is forbidden by symmetry [30], thus V–B emitters intrinsically have low radiative recombination rate and, hence, a low quantum efficiency.
+At the same time, the metastable state has a rather short lifetime (on the order of tens of nanoseconds) [29, 30].
+Therefore, the systems relaxes overall rather rapidly, and the defect can be re-pumped efficiently, such that its PL can still be detected at sufficient intensity despite the intrinsically low quantum efficiency and the presence of the metastable state [30].
 The ODMR spectra are measured at zero external magnetic field and on four different ion fluences (all on the same sample of Figure 1 c) corresponding to four emitter densities (Figure 3 a).
@@
 To extract the splitting between the resonances, we fit the data with two exponentially modified Gaussians, which account for an asymmetric broadening of the resonances (see Figure A.7 b and the discussion in Appendix A.5).
-Figure 3 b shows the resonance positions and , extracted as the maxima of the individual resonance distribution, as a function of the irradiation fluence as well as their midpoint defined as .
+Figure 3 b shows the resonance positions and , extracted as the maxima of the individual resonance distribution, as a function of the irradiation fluence, as well as their midpoint defined as .
 The splitting between the resonances increases with increasing ion fluence, or equivalently defect density (Figure 3 c).
@@
 Following the approach in Refs.
-44, 51, 52, we apply a microscopic charge model to simulate the level splitting.
+44, 54, 55, we apply a microscopic charge model to simulate the level splitting.
 Briefly, we consider level splitting due to a net electric field, which is generated by a discrete and random charge distribution in the vicinity of the V–B site.
@@
 However, unlike Ref.
-52, we do not observe a proportional scaling of charge density and PL intensity with ion fluence, whereby the latter can be taken as a good estimate of the relative number of optically active defects for a constant lifetime (cf.
+55, we do not observe a proportional scaling of charge density and PL intensity with ion fluence, whereby the latter can be taken as a good estimate of the relative number of optically active defects for a constant lifetime (cf.
 Figure 2) and for an excitation power below the saturation threshold (Figure A.5).
 Specifically, the PL intensity of V–B emission increases by a factor of 10 (Figure A.5) as the ion fluence is varied over more than one order of magnitude from to while, at the same time, the average charge density extracted from the model increases only slightly from to .
-Therefore, as a possible extension of the model, we propose that the overall charge density should include a constant positive and negative background charge, which is present before ion irradiation.
-The origin of such a background charge may potentially be related to native defects, such as interstitials, impurities, or vacancies present after the growth.
+Therefore, as a possible extension of the model, we propose that the overall charge density should include a constant positive and negative background charge, potentially due to other defect species.
 Charge transfer processes between such defects can then lead to local charge accumulation, while preserving overall charge neutrality.
@@
 From that, we can find the corrected values for the V–B density as given in Table 1.
+The origin of such a background charge may potentially be related to native defects, such as interstitials, impurities, or vacancies present after the crystal growth and, hence, before the irradiation process.
+Additionally, defects introduced already at very low irradiation fluences, below the rather large fluences needed to create enough optically active V–B sites for reliable ODMR characterization, may contribute substantially to the observed ODMR splitting (see below for a more detailed discussion on the impact of crystal quality).
 III.3 Molecular dynamics simulation of vacancy creation
@@
 Figure 4 b compares the MD approach to simulations based on BCA.
-For the latter we find on average higher defect densities, which agrees with the general notion that these simplified methods tend to overestimate the defect creation.
+For the latter, we find on average higher defect densities, which agrees with the general notion that these simplified methods tend to overestimate the defect creation.
 We thus compare the results from the MD simulations and from the microscopic charge model to estimate the production efficiency of the irradiation process.
 The results are shown in the last column of Table 1.
-Overall, our analysis shows that the production efficiency is rather low at about 0.2 %, implying that although with larger implantation fluences more vacancies are created, a large fraction of vacancies does not get promoted to V–B and remains optically inactive [67, 68, 44].
-It should be pointed out that our simulations do not account for in situ annealing of defects through annihilation of vacancies with interstitials, which are mobile at room temperature [69, 70],as it happens at a macroscopically long time scale.
-The annealing should decrease the number of vacancies, independently of their charge state.
-Our values for the defect yield constitute a lower bound, since we included the potential effect of a constant background charge.
+Overall, our analysis suggests that the production efficiency is rather low at about 0.2 %, implying that, although with larger implantation fluences more vacancies are created, a large fraction of vacancies does not get promoted to V–B and remains optically inactive [71, 72, 44].
+However, the actual fraction of defects that become optically active may be substantially better, as our simulations do not account for in situ annealing of defects through annihilation of vacancies with interstitials, which are mobile at room temperature [73, 74].
+The annealing should decrease the number of created vacancies, independent of their charge state.
+Unfortunately, direct simulation of annealing process using MD is challenging, as annealing happens at macroscopically long time scale, unachievable using MD simulations.
+In addition, charge transfer (e.g., between vacancies and interstitials) and mutual attraction of the defects through screened Coulomb interaction are possible, which should further affect the defect dynamics.
+The only way to model this process is to first perform extensive DFT calculations to assess all migration barriers, accounting for the interaction between the defects, and then perform kinetic Monte Carlo simulations.
+Such simulations will require substantial computational effort and are beyond the scope of this work.
+Overall, our values for the defect yield constitute a lower bound, since we cannot quantitatively account for the impact of defect annealing and since we included the potential effect of a constant background charge in the analysis of the ODMR splitting.
 In contrast, Ref.
@@
 In other words, we tentatively state that the microscopic charge model without background correction tends to significantly overestimate the V–B density at low ion fluences.
+In this context, one may ask how the quality of the hBN crystal will impact the measured OMDR splitting and therefore the corresponding, extracted background charge.
+For this purpose, we compare our ODMR results with previously reported studies that explicitly observed V–B defects and provided sufficient information about the hBN quality, the irradiation parameters, and the ODMR results (Table 3).
+Overall, different studies, using different irradiation methods (both ions and neutrons) and different source crystals, report comparable values of the V–B ODMR splitting (about 100 - 150 MHz).
+Furthermore, a consistent observation throughout is that lower ion fluences result in smaller splitting, in line with our results.
+The lowest values (about 90 MHz) were achieved in high-pressure, high-temperature grown hBN (from NIMS, Japan) [50, 52].
+Indeed, such high-pressure, high-temperature grown hBN is generally considered to be less defective, as evidenced for example by its lower intrinsic defect-related luminescence [75], compared to hBN crystals provided by commercial sources, such as HQgraphene (as used in the present work).
+Yet, we point out that based on the currently available data, we cannot make a conclusively statement on the relation between the hBN quality and the ODMR splitting in the limit of low irradiation fluence.
+Rather a key aspect is that ODMR splittings of irradiation created V–B close to 0 MHz were only reported in few-layer hBN [54].
+Here, it was argued that not the purity of the hBN, but rather the 2D geometry plays a decisive role in reducing the number of nearby background charges resulting in a clean ODMR spectrum with almost negligible peak splitting.
 IV Conclusions
@@
 We analyzed the optical properties of the emitters as a function of ion fluence over more than two orders of magnitude.
-The spatially precise generation allowed us to study the magnetic response probed by ODMR contrast measurements in a consistent and systematic way.
+The spatially precise creation allowed us to study the magnetic response probed by ODMR contrast measurements in a consistent and systematic way.
 We simulated the level splitting obtained experimentally from the ODMR spectra employing a microscopic charge model, where the spatially random distribution of local charges results in a level splitting in ODMR.
@@
 With this modification, we were able to estimate the defect creation yield and compare it to molecular dynamics simulations.
-Our results systematically quantify the He-ion based creation of V–B defects in hBN.
-On the one hand, further developments may involve implementing new theoretical models with the aim of addressing the fundamental question of the origin of the V–B charge [71], such as the nature of the charged state (dynamical or static) and the role of the local defect environment, as has been done for N centers [72].
-On the other hand, spatially precise creation of such emitters as demonstrated here may enable their integration in gated, light-emitting hetero-stacks, in which a tunneling current can be coupled to the in-gap defect states by means of graphene electrodes [56, 73].
+On one hand, further developments may involve the implementation of new theoretical models with the aim of addressing the fundamental question of the origin of the V–B charge [76], such as the nature of the charged state (dynamic or static) and the role of the local defect environment, as it has been done for N centers [77].
+On the other hand, spatially precise creation of such emitters as demonstrated here may enable their integration in gated, light-emitting hetero-stacks, in which the PL emission can be tuned upon application of a bias voltage [78] and a tunneling current can be coupled to the in-gap defect states by means of graphene electrodes [60, 79].
 Acknowledgements.
+We thank Moritz Fischer for helpful discussion and Dorte Rubæk Danielsen for help in the sample fabrication.
 N.
@@
 and I.B.
-acknowledge the NNF Biomag project (NNF21OC0066526).
+acknowledge the NNF Biomag project (NNF21OC0066526) and the Danish National Research Foundation through the center for macroscopic quantum states (bigQ, project no.
+DNRF0142).
 A.V.K.
@@
 Data Availability
-The data that support the findings of this study are available from the corresponding authors upon reasonable request.
+The data that support the findings of this article are openly available [80].
 Appendix A Methods
@@
 A.2 Optical Characterization
+Optical properties are characterized by diffraction-limited confocal scanning photoluminescence (PL) mapping, Raman microscopy, and time-resolved photoluminescence (TRPL) under ambient conditions.
+PL measurements are performed on a custom confocal microscope featuring a 532-nm laser (Cobolt, Hübner GmbH).
+The objective is a Nikon TU Plan Fluor /0.9 NA EPI D.
+Typical laser powers used for the measurements are around 3 mW (0.122 mW measured at the entrance aperture of the objective).
+The signal is collected by a spectrometer (Andor-Solis SR-303i, 150 grooves/mm grating with blaze at 800 nm) equipped with a CCD (Andor Newton).
+Before the He-ion beam irradiation described below, we verify that the selected hBN flakes contain low background luminescence in the spectral range where emission from V–B is expected ( – ) as well as that only a few intrinsic quantum emitters, with typical emission lines in the visible, are present.
+Raman spectroscopy is carried out in a commercial Raman microscope (WiTec Alpha, 532 nm laser, cw-excitation, 300 grooves/mm grating with blaze at 500 nm, Zeiss EC Epiplan-Neofluar objective lens / 0.9 NA).
 | # | (ns) | (ns) |
@@
 | 5 | 1.113(9) | 0.017(3) |
-Optical properties are characterized by diffraction-limited confocal scanning photoluminescence (PL) mapping, Raman microscopy, and time-resolved photoluminescence (TRPL) under ambient conditions.
-PL measurements are performed on a custom confocal microscope featuring a 532-nm laser (Cobolt, Hübner GmbH).
-The objective is a Nikon TU Plan Fluor /0.9 NA EPI D.
-The signal is collected by a spectrometer (Andor-Solis SR-303i, 150 grooves/mm grating with blaze at 800 nm) equipped with a CCD (Andor Newton).
-Before the He-ion beam irradiation described below, we verify that the selected hBN flakes contain low background luminescence in the spectral range where emission from V–B is expected ( – ) as well as that only a few intrinsic quantum emitters, with typical emission lines in the visible, are present.
-Raman spectroscopy is carried out in a commercial Raman microscope (WiTec Alpha, 532 nm laser, cw-excitation, 300 grooves/mm grating with blaze at 500 nm, Zeiss EC Epiplan-Neofluar objective lens / 0.9 NA).
 Time-resolved PL measurements (Figure A.3) are performed on the same setup as for PL experiments.
@@
 The width of the instrument response function (measured on a reflective substrate) is around 250 ps.
-For fitting, we use the Python package Lifefit [74].
+For fitting, we use the Python package Lifefit [81], which uses an FFT-based, iterative reconvolution of exponential decays with an instrument response function to model the lifetime data.
+During fitting, the amplitude weights of the individual decays are optimized using a non-negative least squares algorithm, while the lifetime parameters are optimized by non-linear least squares regression.
+To account for shot noise, Lifefit uses as weights in the least-square optimization, where is the number of photon counts in a time bin.
 A.3 Helium-ion irradiation
-To generate luminescent defects in the hBN films, we use a focused He-ion beam in a helium-ion microscope (Zeiss Orion).
+To create luminescent defects in the hBN films, we use a focused He-ion beam in a helium-ion microscope (Zeiss Orion).
 As irradiation patterns, we choose arrays of circular patches with diameter and 2 µm pitch.
@@
 The interaction of the ion beam with the hBN structure is modeled via an extended Tersoff potential for boron nitride (BN-ExTeP) [83], with a smooth transition to the Ziegler–Biersack–Littmark (ZBL) potential [84] for small distances.
-The simulation spans a material thickness of and is performed for a reference fluence, given by a number of 990 ions and a covered surface of 35 µm 35 µm.
+The simulation spans a material thickness of and is performed for a reference fluence, given by a number of 990 ions and a covered surface of 35 Å 35 Å.
 The MD approach is compared to other Monte Carlo methods, which are the TRIDYN code [85] and the SRIM software [86].
+A.7 Comparison of ODMR splitting across different studies
+In Table 3 we compare our ODMR results with previously reported studies that explicitly observed V–B defects and provided sufficient information about the hBN quality, the irradiation parameters and the ODMR results.
+| Reference | Splitting | hBN source | Irradiation | Fluence |
+| [MHz] | (thickness [nm]) | Method | [× 1015 cm-2] | |
+| This work | HQgraphene (80) | FIB He+ (30 keV) | 1 – 30 | |
+| Gottscholl, Nat. Mater. 2020 [6] | HQgraphene (80000) | Therm. neutrons | 2300 | |
+| Guo, ACS Omega 2022 [21] | HQgraphene (10 - 100) | IB N+ (30 keV) | 0.01 - 1 | |
+| HQgraphene (10 - 100) | IB He+ (30 keV) | 0.1 | ||
+| Ren, J. Lumin. 2023 [23] | HQgraphene (>1000) | FIB He+ (50 keV) | 0.1 | |
+| Gong, Nat. Commun. 2023 [44] | Commerc. avail. hBN | IB He+ (3 keV) | 0.03 - 1 | |
+| Gao, Nano Lett. 2021 [19] | 2D Semiconductors | IB He+ (0.3 - 2.5 keV) | 0.05 | |
+| Zabelotsky, Appl. Nan. Mat. 2023 [51] | 140 | NIMS (100/150) | FIB N+ (12 keV) | 0.625 |
+| Sasaki, Appl. Phys. Lett. 2023 [50] | NIMS (66) | FIB He+ (30 keV) | 1 | |
+| Suzuki, Appl. Phys. Expr. 2023 [52] | 150 | NIMS (>100) | FIB N+ (40 keV) | 1 |
+| 90 (high temp. irr.) | NIMS (>100) | FIB N+ (40 keV) | 1 | |
+| Liang, Adv. Opt. Mat. 2023 [28] | NIMS (>100) | FIB He+ (500 keV) | 1 – 100 | |
+| Udvarhelyi, npj Comput. Mater. 2023 [55] | Own prod. (bulk) | Therm. neutrons | 26 - 260 | |
+| Durand, Phys. Rev. Lett. 2023 [54] | Own prod. (15) | Therm. neutrons | 26 | |
+| Own prod. (few layer) | Therm. neutrons | 26 | ||
+| Haykal, Nat. Commun. 2022 [56] | 120 | Own prod. (bulk) | Therm. neutrons | 26 - 260 |